### v4

# Enhancing Interpretability in Deep Reinforcement Learning through Semantic Clustering

Abstract
In this paper, we explore semantic clustering properties of deep reinforcement learning (DRL) to improve its interpretability and deepen our understanding of its internal semantic organization. In this context, semantic clustering refers to the ability of neural networks to cluster inputs based on their semantic similarity in the internal space. We propose a DRL architecture that incorporates a novel semantic clustering module that combines feature dimensionality reduction with online clustering. This module integrates seamlessly into the DRL training pipeline, addressing the instability of t-SNE and eliminating the need for extensive manual annotation inherent to prior semantic analysis methods. We experimentally validate the effectiveness of the proposed module and demonstrate its ability to reveal semantic clustering properties within DRL. Furthermore, we introduce new analytical methods based on these properties to provide insights into the hierarchical structure of policies and semantic organization within the feature space.
1 Introduction
Deep reinforcement learning (DRL) has been widely applied in domains such as robotics, autonomous systems, game playing, and healthcare, due to its ability to solve complex decision-making tasks [6, 13, 24]. However, the black-box nature of DRL models obscures the decision-making process, potentially leading to unforeseen risks. In this study, we explore semantic clustering properties to improve the interpretability of DRL models. We use the term semantic clustering to refer to the process of grouping information based on underlying semantic similarities in the feature space. Studying semantic clustering in DRL helps reveal the model’s internal knowledge structure and semantic relationships between states, enhancing the interpretability and transparency of DRL models.
Although semantic clustering has been thoroughly investigated in natural language processing (NLP) [21, 20] and computer vision (CV) [14, 22, 19, 34], it remains underexplored in DRL, due to the complexity introduced by temporal dynamics and the absence of direct supervised signals. The sequential nature of decisions in DRL further complicates the task of capturing evolving semantics. Early work introduced external constraints—e.g., bisimulation [9, 33] and contrastive learning [8, 1, 11]—to shape feature spaces conducive to semantic clustering. In contrast, we focus on investigating whether DRL can intrinsically exhibit semantic clustering capabilities.
Mnih et al. [17] and Zahavy et al. [32] analyzed the semantic distribution of the DRL feature space for Atari games using t-distributed Stochastic Neighbor Embedding (t-SNE) [27]. However, these studies are limited in multiple ways: (i) they are limited to a small set of Atari games with fixed scenes, making it difficult to distinguish whether clustering arises from pixel similarity or semantic understanding, (ii) Zahavy et al. [32] manually define features for specific games, imposing substantial human effort, and (iii) both studies rely on t-SNE visualization for semantic analysis, which tends to produce unstable results and lacks an automated clustering mechanism. Thus, these approaches require significant manual effort for feature space annotation and analysis, hindering comprehensive semantic analysis and integration into downstream tasks.
Specifically, we make the following key contributions in this paper:
  • We comprehensively explore the semantic clustering properties of DRL, advancing the understanding of the black-box decision-making processes. Unlike prior work that uses fixed-scene Atari games, we use Procgen  [^1] Detailed environment instructions are available here, in Appendix A of Cobbe et al. [4], and in the repository. [4], which offers rich semantic diversity and dynamic, procedurally generated environments.
  • We introduce a novel architecture that addresses limitations in previous t-SNE-based semantic analysis methods by providing a more stable and effective approach for studying semantic properties in DRL.
  • We present new analysis methods to reveal the internal semantic structure, uncover the hierarchical organization of policies, and identify potential risks in DRL models.
2 Related Work
Semantic Clustering in NLP and CV Prior work in NLP has shown that the spatial arrangement of word embeddings reflects semantic similarities, with semantically-related terms forming clusters in the embedding space [20, 21]. Similarly, in computer vision, images with similar content are positioned closely in the learned feature space [14, 22, 19, 34].
Semantic Clustering in DRL Mnih et al. [17] and Zahavy et al. [32] have previously explored visualizing the DRL feature space using t-SNE. In these studies, t-SNE visualizations show that features of states with close pixel distances tend to cluster together. However, due to the fixed nature of the scenarios they used (Atari games), semantic clustering could not be conclusively verified. This limitation motivates our use of Procgen to validate our approach.
Interpretability of DRL DRL interpretability research often focuses on video games due to their controlled environments and clear rules, which make analyzing decision-making processes easier. PW-Net [10] enhances interpretability by using human-friendly prototypes to explain the model’s decision-making. DIGR [30] improves interpretability by generating saliency maps that highlight the most relevant features influencing the agent’s decisions. Concept policy models integrate expert knowledge into multi-agent RL, enabling real-time intervention and interpretation of agent behavior [31]. MENS-DT-RL [5] applies decision trees to provide a rule-based explanation of the learning process. Furthermore, attention mechanisms and symbolic reasoning frameworks have also been applied to enhance interpretability [25, 18, 15]. Our work explores how DRL models internally organize information, offering new insights into the structure of learned representations.
3 Method
Our proposed architecture with a novel semantic clustering module is presented in Figure 1.
Background
The VQ-VAE [26] workflow begins with an encoder network $\hat{E}$ that maps an input $\mathbf{x}$ to a latent representation $\hat{E}(\mathbf{x})$. This representation is then quantized by mapping it to the nearest embedding in a codebook $\left\{\mathbf{e}_{k}|k\in\left\{1,2,\dots,K\right\}\right\}$. The quantized representation is then passed into a decoder network $\hat{D}$ to reconstruct the input $\mathbf{x}$. The loss function for VQ-VAE is defined as:
$\begin{split}\mathcal{L}_{\text{VQ-VAE}}=\left\|\mathbf{x}-\hat{D}(\mathbf{e}_{k})\right\|_{2}^{2}+\left\|sg\left(\hat{E}(\mathbf{x})\right)-\mathbf{e}_{k}\right\|_{2}^{2}+\beta\left\|sg(\mathbf{e}_{k})-\hat{E}(\mathbf{x})\right\|_{2}^{2}\end{split}$ (1)
where $sg$ is a stop-gradient operator and $\beta$ weights the distance reduction between the encoded output $\hat{E}(\mathbf{x})$ and its closest embedding $\mathbf{e}_{k}$.
In this paper, we modify VQ-VAE to achieve the following objectives: (i) assign features to the nearest VQ embedding for clustering, (ii) seamlessly integrate with DRL training, enabling simultaneous clustering and policy learning, and (iii) enhance clustering and interpretability through joint training with additional losses. Further details are provided in subsection 3.2.
Figure 1: Overview of our architecture. The upper segment represents the classic DRL training pipeline, while the lower segment introduces the semantic clustering module. The Feature Dimensionality Reduction (FDR) net reduces the dimensionality of state features, resulting in FDR features, which the vector quantizer then processes to generate discrete VQ codes (denoted $k$), which represent states associated with clusters, along with the closest VQ embeddings. Subsequently, $k$ is integrated into the state feature by element-wise addition after being expanded to match the state feature dimensions, enabling conditional policy training that better supports the integration of downstream tasks.
3.1 Semantic Clustering Module
To overcome the limitations of previous t-SNE-based semantic analyses (see section 1), we propose a novel semantic clustering module, which includes dimensionality reduction and online clustering.
Dimensionality Reduction Given the complexity of states in most DRL applications, their features are often high-dimensional. For example, Mnih et al. [17] use a 512-dimensional feature vector when training DQN on Atari games. Clustering high-dimensional features is challenging due to "the curse of dimensionality" [3, 2]. To mitigate these issues, we reduce feature dimensionality before clustering, resulting in more robust clustering outcomes. This not only simplifies the clustering process but also enables human-interpretable visualizations, typically in 2D.
The instability of t-SNE arises from its non-convex objective function, making it highly sensitive to initialization and leading to varied and unstable visualization outcomes [27, 28]. To overcome these challenges, we propose the Feature Dimensionality Reduction (FDR) network. This network remaps high-dimensional features to 2D using policy training data for online training, ensuring stable and efficient mappings after training. The FDR network’s loss function is designed to preserve the consistency of distance relationships between high-dimensional and 2D feature spaces, measured by pairwise similarities as described in subsection 3.2.
Online Clustering To reduce the need for extensive manual annotation from previous work and to facilitate the integration of downstream tasks, we propose using online clustering to segment the feature space and assist in semantic analysis. We modify the VQ-VAE framework to achieve online clustering, as detailed in subsection 3.2.
3.2 Loss Function Design
The loss function for our proposed framework is given by
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{DRL}}+f_{\text{control}}\left(w_{\text{FDR}}\mathcal{L}_{\text{FDR}}+w_{\text{VQ-VAE}}\mathcal{L}^{\prime}_{\text{VQ-VAE}}\right).$ (2)
The DRL loss function $\mathcal{L}_{\text{DRL}}$ comes from the original DRL model. $w_{\text{FDR}}$ and $w_{\text{VQ-VAE}}$ are the weights of the FDR loss ($\mathcal{L}_{\text{FDR}}$) and the modified VQ-VAE loss ($\mathcal{L}^{\prime}_{\text{VQ-VAE}}$), respectively. $f_{\text{control}}$ represents the control factor. We explain each of these components below.
FDR Loss $\mathcal{L}_{\text{FDR}}$ is based on state features from the DRL training batch and FDR features generated by the FDR network. We use the Student’s t-distribution for pairwise similarities as it captures nonlinear structures and efficiently measures pairwise relative positions of features within a batch without requiring the entire feature set, making it ideal for online clustering.
The pairwise similarities of state features $p_{ij}$ is given by
$p_{ij}=\frac{d(i,j)}{\sum_{k\neq l}d(k,l)},\quad\text{where }d(m,n)=\left(1+\frac{\left\|f(\mathbf{s}_{m})-f(\mathbf{s}_{n})\right\|^{2}}{\alpha}\right)^{-\frac{\alpha+1}{2}}.$ (3)
where $f$ is the feature extractor, $\mathbf{s}_{i}$ is the $i^{\text{th}}$ state in a batch, and $\alpha$ is the degree of freedom. The pairwise similarities for FDR features, $q_{ij}$ are computed using the same expression as (3), but with $f$ replaced by $g\circ f$, where $g$ is the FDR net. In contrast to other deep clustering studies, e.g., Xie et al. [29] and Li et al. [12], the same degree of freedom $\alpha$ is selected for both high- and low-dimensional similarities, ensuring that the original distance relationship between features is maintained in the low-dimensional space.
The FDR loss is given by
$\begin{split}\mathcal{L}_{\text{FDR}}&=-\sum_{i}\sum_{j}p_{ij}\log\left(q_{ij}\right).\end{split}$ (4)
Modified VQ-VAE Loss To perform clustering, we use the second term of $\mathcal{L}_{\text{VQ-VAE}}$ from (1), which moves VQ embeddings closer to neighboring FDR features. These embeddings function similarly to centroids in online $k$-means [16] clustering. Since the other terms are unnecessary for our model, we only retain and modify the second term to define the modified VQ-VAE loss:
$\mathcal{L}^{\prime}_{\text{VQ-VAE}}=\left\|sg\left[g\left(f\left(\mathbf{s}\right)\right)\right]-\mathbf{e}_{k}\right\|_{2}^{2},$ (5)
where $\mathbf{e}_{k}$ is the closest embedding in the codebook to the FDR feature $g(f(\mathbf{s}))$.
Control Factor Since effective semantic clustering relies on a clear and distinguishable semantic distribution that is often difficult to achieve in the early stages of training, we propose an adaptive control factor ($f_{\text{control}}$) strategy updated according to training performance (see Appendix A).
Improved Clustering Our loss design not only achieves dimensionality reduction and clustering but also enhances clustering properties, making the states within each cluster more compact and the cluster boundaries more separable. This improvement is crucial for clearly distinguishing the semantics of states at the cluster boundaries, further enhancing the model’s interpretability. We demonstrate the improved clustering in section 4 and provide more evidence of this enhanced clustering property and the intrinsic nature of semantic clustering in DRL in subsection C.2.
This enhanced clustering arises from the fact that although the stop-gradient operation ($sg$) in $\mathcal{L}^{\prime}_{\text{VQ-VAE}}$ prevents FDR features from moving towards the closest VQ embeddings, when the FDR features become denser under the influence of $\mathcal{L}_{\text{FDR}}$, it indirectly reduces $\mathcal{L}^{\prime}_{\text{VQ-VAE}}$ and, consequently, $\mathcal{L}_{\text{total}}$. In other words, training with both $\mathcal{L}_{\text{FDR}}$ and $\mathcal{L}^{\prime}_{\text{VQ-VAE}}$ decreases the sum of distances between FDR features and their corresponding VQ embeddings within the same cluster. Additionally, since $\mathcal{L}_{\text{FDR}}$ aims to bring $p$ and $q$ closer, the improved cluster density in the FDR feature space is also transferred to the state feature space.
1
Input: PPO network parameters $\theta$, FDR network parameters $\phi$, SCM hyperparameters, and PPO hyperparameters such as value loss weight $w_{\text{value}}$, entropy loss weight $w_{\text{entropy}}$.
2 for each training iteration $i=1,2,\dots$ do
  Collect NN trajectories $\mathcal{D}_{i}=\{\tau_{1},\dots,\tau_{N}\}$ using policy $\pi_{\theta}$;
 // Trajectory collection
3 for each epoch $j=1,2,\dots$ do
4    for each minibatch $M\subseteq\mathcal{D}_{i}$ do
5       for each state $\mathbf{s}_{m}\in M$ do
          $\mathbf{f}_{m}\leftarrow f_{\theta}(\mathbf{s}_{m})$;
          // Extract state feature
          $\mathbf{f}_{m}^{\text{FDR}}\leftarrow g_{\phi}(\mathbf{f}_{m})$;
          // Extract FDR feature
          $k_{m}\leftarrow\arg\min_{k}\|\mathbf{f}_{m}^{\text{FDR}}-\mathbf{e}_{k}\|$;
          // Assign to nearest VQ embedding
          $\mathbf{k}_{m}^{\text{expand}}\leftarrow\text{expand}(k_{m},\dim(\mathbf{f}_{m}))$;
          // Broadcast to state feature dim.
          $\mathbf{f}_{m}^{\text{fused}}\leftarrow\mathbf{f}_{m}+\mathbf{k}_{m}^{\text{expand}}$;
          // Apply element-wise addition
          $\pi(a|\mathbf{s}_{m})\leftarrow\hat{\pi}_{\theta}(\mathbf{f}_{m}^{\text{fused}})$;
          // Compute policy outputs
          $V(\mathbf{s}_{m})\leftarrow\hat{V}_{\theta}(\mathbf{f}_{m}^{\text{fused}})$;
          // Compute value outputs
6
       $\mathcal{L}_{\text{PPO}}\leftarrow\mathcal{L}_{\text{policy}}+w_{\text{value}}\cdot\mathcal{L}_{\text{value}}+w_{\text{entropy}}\cdot\mathcal{L}_{\text{entropy}}$;
       // PPO loss
       $\mathcal{L}_{\text{SCM}}\leftarrow w_{\text{FDR}}\cdot\mathcal{L}_{\text{FDR}}+w_{\text{VQ-VAE}}\cdot\mathcal{L}^{\prime}_{\text{VQ-VAE}}$;
       // SCM loss
       $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{PPO}}+f_{\text{control}}\cdot\mathcal{L}_{\text{SCM}}$;
       // Total loss
         Update $\theta$, $\phi$, and $\{\mathbf{e}_{k}\}_{k=1}^{K}$ by minimizing $\mathcal{L}_{\text{total}}$;
       // Parameter update
7
8
9
Algorithm 1 PPO with Semantic Clustering Module (SCM)
Advantages of Online Training Online training offers several key advantages: (i) It enhances clustering by incorporating the training of $\mathcal{L}_{\text{total}}$. (ii) Training the VQ code $k$ with a latent-conditioned policy $\pi\left(a|\mathbf{s},k\right)$ (where $a$ is the action) supports the extension to downstream tasks, such as macro action selection in hierarchical learning. (iii) Online training improves memory efficiency by eliminating the need to store a large number of states during model training.
Training Process The training process of our framework builds upon the structure of the original DRL algorithm while incorporating the semantic clustering module (SCM) by using (2) for total loss calculation. We take PPO [23] as an example, and the training procedure is outlined in algorithm 1.
4 Simulations
In this work, we primarily study the intrinsic characteristics and black-box decision-making of DRL, and address the instability of t-SNE visualizations used in prior studies. Therefore, this section aims to: (i) compare t-SNE to validate the stability and effectiveness of the proposed clustering method; (ii) assess the semantic clustering properties of DRL to improve interpretability; and (iii) introduce new methods to analyze policies and internal model characteristics, identifying issues in DRL decision-making. For performance analysis, please refer to subsection C.1 and Appendix D, which demonstrate that the integration of our module has minimal impact on performance.
4.1 Clustering Effectiveness Evaluation
(a) t-SNE space of PPO.
(b) t-SNE space of our method.
(c) Same as (b), but with 50% fewer states.
(e) Same as (d), but with 50% fewer states.
(f) Same as (c), but with a different random seed.
Figure 2: Visualization of features in t-SNE and FDR spaces using PPO and our method. To enable comparison, feature colors in the t-SNE visualizations of our method correspond to the cluster colors in the FDR space, while PPO features are shown in orange due to the absence of clustering. Unlike t-SNE, which fails to produce clearly separable clusters and exhibits sensitivity to the number of states and random seeds, our method yields well-separated and stable clusters under varying conditions.
(d) FDR space of our method.
We demonstrate the clustering effectiveness of our proposed approach using the CoinRun game from Procgen as an example. Similar results can be easily extended to other games using the code and checkpoints provided in the supplementary material. We use a trained model to collect states, where the agent selects actions randomly with a probability of 0.2 and follows the trained policy with a probability of 0.8 to ensure diverse state coverage. States are sampled with a probability of 0.8, and 64 parallel environments collect states over 500 steps, resulting in $\sim$25,000 states for visualization.
Cluster Separation and Improved Clustering The t-SNE visualization of PPO (2(a)), spreads features across the space without forming clear clusters, limiting its utility for clustering analysis and requiring detailed manual examination of certain areas, as in previous studies. In contrast, the t-SNE visualization of our method (2(b)), reveals numerous distinct, small clusters. States within each of these clusters originate from the same semantic group identified by our method. This dispersion into multiple smaller clusters is due to t-SNE’s focus on local structures and its tendency to avoid crowding, causing complete semantic clusters to scatter. The visualization in the FDR space (2(d)), displays clear and separate complete clusters, which are identified by VQ codes.
Sensitivity to Number of States Our method’s stability is showcased in Figures 2(c) and 2(e), where the number of processed states is reduced by 50%. Unlike the drastic changes in feature distribution seen in the t-SNE space (2(c)), the FDR space (2(e)) exhibits a stable mapping, merely reducing the quantity of features without altering their spatial distribution.
Sensitivity to Random Seed While the t-SNE representation is sensitive to randomness, as demonstrated by the significant difference between Figures 2(c) and 2(f), the FDR space’s mapping remains unchanged even when the random seed is altered, maintaining the distribution in 2(e). t-SNE’s randomness primarily stems from its random initialization and non-convex optimization process, leading to different visualizations with different random seeds. In contrast, our model produces a stable feature mapping after training, which does not vary with random seeds.
These clear contrasts highlight the robustness of FDR over the instability of t-SNE, addressing prior limitations and enabling stable semantic clustering and analysis. In Appendix E, we further present a statistical comparison of common dimensionality reduction methods across multiple clustering metrics, demonstrating that our method achieves superior clustering performance.
4.2 Semantic Clustering in DRL
Figure 3: State examples in the Ninja FDR space and the mean images of clusters. Each dashed box contains a sequence of consecutive states assigned to the same cluster, with dotted arrows indicating their corresponding FDR feature positions. These examples demonstrate that semantically similar and temporally adjacent states are grouped into the same cluster, highlighting the learned semantic coherence. Descriptions of the state sequences in the clusters are provided in Table 1.
Table 1: Cluster descriptions and mean image outlines for the Ninja game
Cluster Description Mean image outlines
0 The agent starts by walking through the first platform and then performs a high jump to reach a higher ledge. Essential elements are outlined, e.g., a left-side wall, the current position of the agent on the first platform, and the upcoming higher ledges.
1 The agent makes small jumps in the middle of the scene. We can observe the outlines of several ledges below the agent.
2 Two interpretations are present: 1) the agent starts from the leftmost end of the scene and walks to the starting position of Cluster 0, and 2) when there are no higher ledges to jump to, the agent begins from the scene, walks over the first platform, and prepares to jump to the subsequent ledge. The scene prominently displays the distinct outline of the left wall and the first platform. The agent’s current position is close to both of them.
3 The agent walks on the ledge and prepares to jump to a higher ledge. The agent is standing on the outline of the current ledge and the following higher ledges.
4 After performing a high jump, the agent loses sight of the ledge below. The agent is performing a high jump.
5 The agent walks on the ledge and prepares to jump onto a ledge at the same height or lower. The agent is standing on the outline of the current ledge and the following ledges at the same height or lower.
6 The agent executes a high jump while keeping the ledge below in sight. The agent is performing a high jump and the outline of the ledge below is visible.
7 The agent moves towards the right edge of the scene and touches the mushroom. The outlines of the wall and platform on the far right are visible.
In this section, we illustrate semantic clustering analysis using the Ninja game, in which the agent goes from left to right, jumping over various ledges and scores points by touching the mushroom on the far right. In Appendix G, we analyze additional games, reaching similar conclusions.
Mean Image Analysis We performed a qualitative analysis of the mean images of states within each cluster. Figure 3 presents state examples from the FDR space of Ninja along with the mean images of each semantic cluster, and Table 1 contains natural language descriptions of the clusters as well as notable features of the mean images corresponding to each cluster. Corresponding videos can be found in the supplementary material.
Unlike static semantic clustering in some CV and NLP tasks, where clustering is based on a single image or word, DRL’s semantic clustering is dynamic in nature—state sequences with similar semantics are grouped into the same semantic cluster. Notably, this semantic clustering goes beyond pixel distances and operates on a semantic understanding level of the environment, as illustrated in figures 3 and 4. This generalized semantic clustering emerges from the DRL model’s inherent ability to learn and summarize from changing scene dynamics, independent of external constraints like bisimulation or contrastive learning, and without the need for supervised signals. The neural network’s internal organization of policy-relevant knowledge indicates clustering-based spatial organization based on semantic similarity. Furthermore, we find that video sequences within clusters can be summarized using natural language, akin to the ‘skills’ humans abstract during learning processes.
Figure 4: Three episodes from the Ninja game. States within colored dashed boxes correspond to clusters of the same colors in Figure 3. Solid gray arrows indicate omitted intermediate states from the same cluster, while ellipses represent other omitted states. These visualizations illustrate consistent semantic alignment in cluster assignments across different episodes.
Table 2: Human evaluation results
Mean Score (SEM)
No. Statement Jumper FruitBot Ninja
1 The clips of each cluster consistently display the same skill being performed 4.24 (0.15) 4.10 (0.11) 4.30 (0.15)
2 The clips of each cluster match the given skill description 4.36 (0.16) 4.16 (0.11) 4.20 (0.17)
3 The identified skills aid in understanding the environment and the AI’s decision-making process 4.50 (0.22) 4.10 (0.18) 4.20 (0.20)
Human Evaluation In addition to qualitatively analyzing the mean images, we hired 15 human evaluators to validate the semantic clustering properties. Specifically, video sequences from each episode are segmented into multiple clips based on the cluster each frame belongs to, and these clips are grouped by cluster for evaluators to review. Each evaluator watched these grouped clips and responded to three interpretability-related statements for two out of a set of three games (Jumper, Fruitbot, and Ninja). The response for each question was chosen from a five-point Likert scale with the following options: Strongly Disagree (1), Disagree (2), Neutral (3), Agree (4), and Strongly Agree (5). Further details on the human evaluation procedure are provided in Appendix I.
The statements and the results of the human evaluation are provided in subsection 4.2. The mean scores for all statement-environment combinations are greater than 4, with the exception of statements 1 and 3 for FruitBot, for which the lower bounds on the mean set by the standard error of the mean (SEM) are 3.99 and 3.92 respectively. The slightly lower score on FruitBot may be caused by the behavior description of clusters, the agent’s relative distance to the wall ahead (far/near) and the agent’s relative position on the screen (left/center/right) require a higher degree of subjective judgment. In contrast, Jumper has a clear radar for direction and position information, and Ninja has more explicit behavioral reference objects, e.g., ledges and mushrooms. Overall, these results suggest that humans generally agree that our model possesses semantic clustering properties and supports interpretability.
4.3 Model and Policy Analysis
(a) Overall perspective.
(b) First example point
(c) Second example point
(d) Third example point
Figure 5: Hover examples in the FDR space of Ninja. We observe a sub-cluster in the FDR space as an example from the overall perspective (a) and the zoomed-in perspective (b), (c), and (d). The agent is standing on the edge of a ledge. Although the scenarios of (b), (c), and (d) are different, the proposed method effectively clusters semantically consistent features together in the FDR space.
To better explore the knowledge organization within the internal space of DRL models, we developed a visualization tool (see Figure 5 for an example). The tool supports ‘statically’ analyzing the semantic distribution of models—specifically, (i) when the mouse cursor hovers over a specific feature point, the corresponding state image is displayed, and (ii) the tool includes a zooming functionality to observe the semantic distribution of features in detail within clusters.
Figure 6: Policy analysis examples in Ninja.
In addition, we propose a more ‘dynamic’ analysis method—the VQ code enables us to determine the cluster to which the current state belongs, which allows for the semantic segmentation of episodes, as exemplified in Figure 4. Our model excels at breaking down complex policies, thereby shedding light on their inherent hierarchical structures. Moreover, this segmentation is based on semantics, making it understandable to humans and likely to improve interpretability in downstream hierarchical learning tasks. Consequently, this method introduces a ‘dynamic’ strategy for dissecting policy structures.
We present policy analysis examples in Figure 6, leveraging clustering results from our method. Figures 6(a) and 6(b) illustrate consecutive states assigned to clusters 5 and 7, respectively. These examples show that the decision to approach the mushroom depends not on the mushroom’s presence but on detecting a right-side wall. This finding is confirmed in 6(c), another state from cluster 7 without any mushroom present, where the agent repeatedly moves along the ledge, incorrectly perceiving conditions for behaviors in cluster 7. Figures 6(d) and 6(e) depict states within the same episode: in 6(d), the agent initially plans to jump onto a lower ledge (cluster 5), but upon seeing a higher, safer ledge in 6(e), it shifts its strategy accordingly (cluster 3). Similarly, 6(f) shows a state that has been assigned to cluster 3 by our model, helping us anticipate the agent’s future behavior of jumping onto the higher ledge. These analyses demonstrate how our model helps clarify policy behaviors, uncover decision-making structures, and identify potential issues.
5 Limitations and Future Work
Our approach has several limitations to be addressed in future work. First, it relies on clear semantic distributions, which can become unstable when policies deviate significantly from optimal behavior, resulting in ambiguous clusters. More robust clustering methods may be needed to improve stability. Second, as the method is unsupervised, selecting an appropriate number of clusters is crucial—too few reduce clarity, while too many cause semantic fragmentation. We used eight clusters to balance interpretability and granularity, but future work could explore adaptive techniques that adjust the number of clusters based on task complexity. For further analysis of the impact of cluster numbers on performance and interpretability, see Appendix D. Furthermore, policy interpretations are manually described. In future work, we aim to automate behavior summarization and explanation. Lastly, we plan to extend this method to other DRL algorithms, benchmarks, and settings.
6 Conclusion
In this paper, we investigated the semantic clustering properties of DRL. Using a novel approach that combines dimensionality reduction and online clustering, we analyzed the internal organization of knowledge within the feature space. Our method provides a stable mapping of feature positions and enhances semantic clustering, revealing meaningful structures in continuous sequences of video game states. We demonstrate that semantic clustering in DRL arises dynamically as the agent interacts with its environment. As the agent explores diverse states during reinforcement learning, it naturally clusters semantically related states based on spatial and temporal relationships. This dynamic clustering exploits regularities in the environment, offering a unique approach compared to the static clustering observations in NLP and CV.
Appendix Overview
Appendix A Architecture, Hyperparameters, and Computational Costs
The training of the proposed method is consistent with the Impala architecture [7], the PPO algorithm [23], and the hyperparameters used in the Procgen paper [4]. The FDR net is composed of two fully connected layers with 128 and 2 neurons, respectively. The codebook in the vector quantizer has eight embeddings, and the degree of freedom in the FDR loss is 20. The random seeds employed in 2(c) are 2021 and 2031, while the seeds used in Figure C.1 are 2021, 2022, and 2023. We train all models on one NVIDIA Tesla V100S 32GB GPU. The operating system version is CentOS Linux release 7.9.2009. Each runs takes around six hours.
In Equation 2 of the main paper, $w_{\text{FDR}}$ and $w_{\text{VQ-VAE}}$ are 500 and 1, respectively. $f_{\text{control}}$ is updated every 50 iterations according to the following expression:
$f_{\text{control}}=\min\left(\frac{s_{\text{mean}}}{0.8\cdot s_{\text{highest}}},1\right),$ (A.1)
where $s_{\text{mean}}$ is the mean score of the last 100 episodes in training, and $s_{\text{highest}}$ is the highest score of the environment.
All hyperparameters introduced in our method, except for the number of embeddings, were chosen through performance tuning to optimize the model’s overall performance. The number of embeddings in the vector quantizer was determined by ensuring that each cluster maintained a singular semantic interpretation.
Appendix B Theoretical Analysis of Loss Design
The two auxiliary losses in our framework serve distinct theoretical purposes. The $\mathcal{L}_{\mathrm{FDR}}$ term is intended to preserve relative geometry when mapping high‑dimensional state features to a 2‑D space, whereas the modified VQ term should behave like a standard online kk‑means step so that the codebook converges to meaningful cluster centroids. We formalize both claims below.
B.1 Batch-wise Distance Similarity Preservation
Goal.
We first prove that driving $\mathcal{L}_{\mathrm{FDR}}\to 0$ guarantees that pairwise similarity orderings are preserved between the original feature space and the FDR space, thereby retaining semantic neighborhood structure.
Notation.
Let
  • $n$ be the mini‐batch size.
  • $\{x_{i}\}_{i=1}^{n}$ be the batch of high‐dimensional state features.
  • $\{y_{i}\}_{i=1}^{n}$ be the corresponding low‑dimensional embeddings produced by the FDR network.
  • $\alpha>0$ be the degrees of freedom of the Student–$t$ kernel.
  • The Student–$t$ kernel
    $d_{\mathrm{t}}(u,v)=\Bigl(1+\tfrac{\|u-v\|^{2}}{\alpha}\Bigr)^{-\frac{\alpha+1}{2}},\quad d_{\mathrm{t}}(u,v)\in(0,1].$
Define for all $1\leq i\neq j\leq n$ the normalized pairwise similarities
$p_{ij}=\frac{d_{\mathrm{t}}(x_{i},x_{j})}{\sum_{k\neq\ell}d_{\mathrm{t}}(x_{k},x_{\ell})},\qquad q_{ij}=\frac{d_{\mathrm{t}}(y_{i},y_{j})}{\sum_{k\neq\ell}d_{\mathrm{t}}(y_{k},y_{\ell})},$
and the Kullback–Leibler divergence
$\mathcal{L}_{\mathrm{FDR}}=\sum_{i\neq j}p_{ij}\,\log\frac{p_{ij}}{q_{ij}}\;\geq\;0.$
Theorem 1 (Similarity Preservation).
If $\mathcal{L}_{\mathrm{FDR}}=0$, then there exists a constant
$\kappa\;=\;\frac{\sum_{k\neq\ell}d_{\mathrm{t}}(y_{k},y_{\ell})}{\sum_{k\neq\ell}d_{\mathrm{t}}(x_{k},x_{\ell})}\;>\;0$
such that for all $i\neq j$,
$d_{\mathrm{t}}(y_{i},y_{j})\;=\;\kappa\,d_{\mathrm{t}}(x_{i},x_{j}).$
Consequently, the ordering of squared distances $\|y_{i}-y_{j}\|^{2}$ and $\|x_{i}-x_{j}\|^{2}$ is identical. Moreover, if $\kappa=1$, then $d_{\mathrm{t}}(y_{i},y_{j})=d_{\mathrm{t}}(x_{i},x_{j})$ and hence $\|y_{i}-y_{j}\|^{2}=\|x_{i}-x_{j}\|^{2}$.
Proof.
By the non‑negativity of KL divergence, $\mathcal{L}_{\mathrm{FDR}}=0$ iff $p_{ij}=q_{ij}$ for every $i\neq j$. Equating
$\frac{d_{\mathrm{t}}(y_{i},y_{j})}{\sum_{k\neq\ell}d_{\mathrm{t}}(y_{k},y_{\ell})}\;=\;\frac{d_{\mathrm{t}}(x_{i},x_{j})}{\sum_{k\neq\ell}d_{\mathrm{t}}(x_{k},x_{\ell})}\quad\Longrightarrow\quad d_{\mathrm{t}}(y_{i},y_{j})=\kappa\,d_{\mathrm{t}}(x_{i},x_{j}).$
Since the Student‑$t$ kernel decreases strictly with $\|u-v\|^{2}$, scaling by $\kappa>0$ preserves rank order. When $\kappa=1$, strict monotonicity forces equality of squared distances. ∎
Minimizing $\mathcal{L}_{\mathrm{FDR}}$ therefore enforces a batch‑wise isometry up to a global scale factor $\kappa$, which is exactly the property needed for semantic clustering in the 2‑D FDR space.
B.2 Modified VQ–kk‐Means Equivalence
Goal.
Next we show that the gradient update used for the vector‐quantizer codebook is algebraically identical to an online kk‑means step.
Notation.
Let
  • $y_{i}=g(f(x_{i}))$ be the FDR feature for state $x_{i}$.
  • $\{e_{k}\}_{k=1}^{K}$ be the learnable codebook entries.
  • $k_{i}=\arg\min_{1\leq k\leq K}\|y_{i}-e_{k}\|_{2}$ be the nearest code index.
  • $m_{k}=\lvert\{i:k_{i}=k\}\rvert$ be the cluster size.
  • $\bar{y}_{k}=\tfrac{1}{m_{k}}\sum_{i:k_{i}=k}y_{i}$ be the empirical cluster centroid.
  • The modified VQ loss $\mathcal{L}^{\prime}_{\mathrm{VQ}}=\sum_{i=1}^{n}\bigl\|\mathrm{sg}(y_{i})-e_{k_{i}}\bigr\|_{2}^{2}.$
Theorem 2 (Equivalence to Online kk‑Means).
Updating each codebook vector via
$e_{k}^{(t+1)}=e_{k}^{(t)}-\beta\,\nabla_{e_{k}}\mathcal{L}^{\prime}_{\mathrm{VQ}},\qquad\beta>0,$
produces
$e_{k}^{(t+1)}=e_{k}^{(t)}+\gamma_{k}\bigl(\bar{y}_{k}-e_{k}^{(t)}\bigr),\qquad\gamma_{k}=2\beta\,m_{k},$
which is the online kk‑means update with learning rate $\gamma_{k}$.
Proof.
Because stop‑gradient $\mathrm{sg}(y_{i})$ blocks gradients with respect to $y_{i}$,
$\nabla_{e_{k}}\mathcal{L}^{\prime}_{\mathrm{VQ}}=2\sum_{i:k_{i}=k}(e_{k}-y_{i}).$
Therefore
$e_{k}^{(t+1)}=e_{k}^{(t)}-2\beta\sum_{i:k_{i}=k}\bigl(e_{k}^{(t)}-y_{i}\bigr)=e_{k}^{(t)}+2\beta\,m_{k}\bigl(\bar{y}_{k}-e_{k}^{(t)}\bigr),$
which matches the standard incremental kk‑means rule. ∎
Appendix C Ablation Study on Performance and Interpretability
C.1 Performance Impact of the Semantic Clustering Module
Considering the cost of time and computational resources, we opt for training our model on the full distribution of levels in the ‘easy’ mode. In Figure C.1, a comparison of performance curves between the proposed method and the baseline is presented, where ‘SPPO’ denotes ‘semantic’ PPO, i.e., PPO integrated with our proposed semantic clustering module. Consistent with the Procgen paper [4], given the diversity of episodes during training, a single curve represents both training and testing performance. Across these environments we observe that the proposed method closely aligns with the baseline performance, indicating minimal impact on performance from the introduced module. This is expected, as the module only performs dimensionality reduction and clustering based on existing features, without introducing external information. The discrete code $k$ reflects only the position of a feature in the learned space and is expanded and added element-wise to preserve the original feature dimensionality, ensuring the policy receives no additional information beyond what is already contained in the state feature.
Figure C.1: Performance curves on ‘easy’ difficulty environments using three random seeds, trained and evaluated on the full distribution of levels.
C.2 Semantic Clustering as an Intrinsic Property of DRL
We conducted a stop gradient experiment to further investigate whether semantic clustering is an inherent property of DRL. In this experiment, we applied a stop gradient operation to Equation 4 and removed the connection between the VQ codes and the original state features. This was done to prevent the semantic clustering module from influencing the feature space and to observe whether semantic clustering would still occur. The results, as shown in Figures C.2 and C.3, demonstrated that states within the same semantic cluster continued to exhibit similar semantic interpretations, even without the influence of the semantic clustering module. However, the boundaries between clusters became less clear, making it more difficult to distinguish the semantics of states near the edges of clusters. Notably, we also observed that this modification had minimal impact on performance, consistent with the trend shown in Figure C.1.
Figure C.2: Visualization of Features in the t-SNE Space. The training eliminates the impact of the proposed semantic clustering module on the original feature space. Feature colors correspond to cluster colors in the FDR space of Figure C.3, facilitating the comparison of spatial relationships and feature distribution changes. Compared to 2(b) in the main paper, the absence of the semantic clustering module’s enhancement makes sub-clusters less distinct.
Figure C.3: Visualization of Features in the FDR Space. The training eliminates the impact of the proposed semantic clustering module on the original feature space. Compared to 2(d) in the main paper, the cluster boundaries in the FDR space are less distinct.
These observations suggest that semantic clustering is indeed an intrinsic property of DRL, driven by the agent’s interaction with its environment during training. The proposed semantic clustering module enhances this natural clustering behavior by increasing the density of clusters, thus improving the separability between them. To fine-tune the influence of the module, we introduced a control factor. At the beginning of training, the control factor is kept low, allowing the DRL training to shape the feature space independently. As the policy becomes more optimized and the semantic distribution of states becomes more organized, the control factor is gradually increased to further enhance the clarity and separability of clusters.
Appendix D Impact of the Number of VQ Embeddings on Performance and Interpretability
To analyze the effect of the number of VQ embeddings ($K$) on both model performance and interpretability, we conducted experiments using the Jumper environment as an example. Similar conclusions can be extended to other environments.
D.1 Performance Analysis
Figure D.1 shows the performance of our model with varying numbers of VQ embeddings. The results demonstrate that the number of embeddings does not affect model performance. This is expected, as our proposed method primarily focuses on feature dimensionality reduction and clustering. Combined with the overall performance results in Figure C.1, we observe that the model maintains consistent performance.
Figure D.1: Performance comparison of models with different numbers of VQ embeddings in the Jumper environment.
D.2 Interpretability Analysis
Figures 2(a) and 2(b) illustrate the FDR results for $K=4$ and $K=12$, respectively. Our method effectively produces clusters that are clearly separable, regardless of the VQ embedding number.
However, interpretability is influenced by the choice of $K$. When $K=12$, the semantic clusters become overly fragmented, making it difficult to form coherent semantic explanations for clusters. Conversely, when $K=4$, Table 3 shows that clusters contain multiple distinct semantic explanations, which negatively impacts interpretability.
(a) $K=4$
(b) $K=12$
Figure D.2: Visualization of the FDR spaces for different numbers of VQ embeddings in the Jumper environment.
Table 3: Cluster descriptions for the Jumper game with $K=4$
Cluster Description
0 1) The agent is touching the carrot on the upper left.
2) The agent is touching the carrot on the right.
3) The agent is touching the carrot on the bottom right.
4) The agent is moving in the left or lower-left part of the scene.
1 1) The agent is touching the carrot above.
2) The agent is touching the carrot on the left.
3) The agent is moving in the right or lower-right part of the scene.
2 1) The agent is touching the carrot below.
2) The agent is moving in the upper part of the scene.
3 1) The agent is moving at the bottom of the scene.
2) The agent is approaching the carrot above.
Clusters with incomplete or incoherent semantic descriptions hinder interpretability by introducing ambiguity in understanding the agent’s behavior. This lack of clarity complicates policy analysis and makes it challenging to draw meaningful insights. Conversely, when a single cluster contains multiple interpretable behaviors, it increases the cognitive load for users who must disambiguate between these behaviors. Such a many-to-one mapping between behaviors and clusters undermines the straightforward identification of the agent’s current strategy, reducing the utility of clustering as a tool for decision-making. To address these challenges, it is essential to ensure a one-to-one mapping between clusters and explanations. When each cluster is associated with a single, coherent explanation, it eliminates the need for further distinctions within clusters, facilitating clear policy analysis and enhancing human understanding of the agent’s behavior.
Appendix E Clustering Quality under Different Dimensionality‑Reduction Methods
(a) Silhouette score (higher is better)
(b) Davies–Bouldin Index (lower is better)
(c) Calinski–Harabasz score (higher is better)
Figure E.1: Clustering metrics averaged over three Procgen games (Ninja, Jumper, and Fruitbot), each run with three random seeds (nine runs in total). Bars show mean values; error bars denote the standard error of the mean.
To quantify how well various dimensionality‑reduction (DR) techniques preserve the semantic clusters discovered by our model, we project the same set of high‑dimensional state features—collected following the procedure described in subsection 4.1—using five popular DR baselines: UMAP, Isomap, PCA, MDS, and t‑SNE. These metrics were selected to evaluate both the compactness within clusters and the separation between clusters in the reduced feature space.
  • Silhouette – cohesion vs. separation of each point’s cluster.
  • Davies–Bouldin Index (DBI) – average cluster similarity (lower indicates tighter, more separated clusters).
  • Calinski–Harabasz (CH) – ratio of between‑ to within‑cluster dispersion.
Figure E.1 shows that our learned 2‑D FDR space achieves the best scores on all metrics: Silhouette is the highest, DBI the lowest, and CH an order of magnitude larger than any baseline. These results confirm that our dimensionality‑reduction module preserves—and even sharpens—the intrinsic semantic clustering properties uncovered in the high‑dimensional feature space.
Appendix F Semantic Formation in Clusters
To provide insight into the generation of semantic clustering in the feature space, we conducted the statistical analysis as follows. Initially, we collected 50,000 states using the trained model and computed the mean image for each cluster. We then calculated the mean and standard deviation of the pixel distances between the states in each cluster and their corresponding mean image. These measurements were averaged across the eight clusters. Additionally, we evaluated the probability of cluster transitions in these 50,000 states. We repeated the same statistical analysis on the 50,000 states using the untrained model, and the results are presented in Table 4. In the trained models, the pixel differences between the states within clusters increased, indicating the presence of states with similar semantics but larger pixel distances. Furthermore, the probability of cluster transitions decreased as it approached the required transition in the scene.
Table 4: Comparison metrics of the semantically trained model verse the raw model on various environments.
Model Cluster transition probability Pixel distance mean (Std. Dev.)
FruitBot Trained 0.1081 100.00 (71.33)
FruitBot Raw 0.2834 77.10 (49.94)
Jumper Trained 0.2224 110.29 (62.29)
Jumper Raw 0.5829 104.57 (58.31)
Ninja Trained 0.2680 141.50 (67.88)
Ninja Raw 0.2712 87.61 (62.43)
Appendix G More Examples and Mean Images in the FDR Space
G.1 CoinRun
To augment the exploration of semantic clustering as discussed in the main paper, this section analyzes two additional games characterized by distinct dynamics. CoinRun’s gameplay mechanism is similar to Ninja’s, requiring the agent to traverse from the far-left to the far-right, scoring points by interacting with coins at the far-right end of the scene, as illustrated in Figure G.1.
Figure G.1: A episode in CoinRun. Ellipses represent the omitted states.
The observations and insights obtained closely mirror those derived from the analysis of Ninja. Interested readers can leverage the provided code and checkpoint for further exploration of similar findings. Consequently, for brevity, we refrain from extensively elaborating on analogous conclusions.
G.2 Jumper
In Jumper, the agent navigates a cave to locate and touch carrots by interpreting a radar displayed in the upper right corner of the screen. The radar’s pointer indicates the direction of the carrot, while a bar below the radar shows the distance between the agent and the carrot—shorter bars imply closer proximity, and vice-versa.
Figure G.2: Examples and mean images from the Jumper FDR space.
Table 5: Cluster descriptions and mean image outlines for the Jumper game
Cluster Description Mean image outlines
0 The agent learns to jump up from the bottom left and move to the left on the top right. The radar pointing up and to the right, and the outline of the channel above and to the right..
1 The agent is touching the carrot on the left or upper left. The radar is pointing to the upper left and is very close to the target.
2 The agent learns the skill of movement at the top of the scene. The radar pointing to the left or down, and the outline of the channels face to the left or down.
3 The agent is approaching the carrot on the upper right. The radar pointing to the upper right, and the distance to the target is very close.
4 The agent is touching the carrot below. The radar is pointing down, and it is very close to the target.
5 The agent is approaching the carrot above or left. The radar is pointing up, and it is very close to the target.
6 The agent is touching the carrot on the right. The radar is pointing right, and it is very close to the target.
7 The agent learns the skill of movement at the right bottom of the scene. The radar pointing up or to the top left, and it is far from the target.
The state examples and mean images from the clusters in the FDR space of Jumper are presented in Figure G.2. The background of Jumper is diverse, and the agent is always in the center of the screen (zoom in to see the outlines clearly). In Table 5, we break down descriptions of the sampled images from each cluster and interpretations of the mean image for each cluster in the Jumper game.
Figure G.3 depicts various states from the Jumper game. C.3(a) and C.3(b) belong to the same episode and fall under Cluster 4, while C.3(c) and C.3(d) are from another episode, both categorized under Cluster 1. Notably, neither C.3(a) nor C.3(c) shows the presence of carrots. This observation leads us to suspect that the determination of these clusters is solely reliant on the radar and distance bar rather than the appearance of carrots. To test this hypothesis, we removed the carrot in C.3(e), which originally belonged to Cluster 4, and transformed it into C.3(f). The result demonstrated that C.3(f) still belongs to Cluster 4, confirming our suspicion. However, this phenomenon might pose potential risks in practical applications. For example, in scenarios where sensor data and visual perceptions misalign, AI models might solely rely on sensor data for decision-making—e.g., an autonomous vehicle’s sensors indicating an empty road while the occupants inside observe pedestrians crossing, yet the vehicle continues to accelerate.
Figure G.3: Policy analysis examples in Jumper.
G.3 FruitBot
FruitBot is a bottom-to-top scrolling game where the agent moves left or right to collect fruits for points while avoiding negative scores upon touching non-fruit objects. The state examples and mean images from the clusters in the FDR space of FruitBot are presented in Figure G.4 and their descriptions in Table 6. FruitBot’s mean images lack clarity due to the presence of diverse backgrounds, and the agent is constantly moving to the left and right at the bottom of the screen. However, we can still make out the outline of the wall and agent if we look carefully (zoom in to see the outlines clearly).
Figure G.4: State examples and mean images from the FruitBot FDR space.
Table 6: Cluster descriptions and mean image outlines for the FruitBot game
Cluster Description Mean image outlines
0 The agent is approaching the wall in the left area. We can see the agent moving toward the gap on the wall that is approaching on the lower left.
1 The agent approaches the wall from the right area. The agent is moving toward the gap on the wall that is approaching on the lower right.
2 The agent executes its policy far from the wall from the left area. The wall is far away, and the agent is moving in the lower left.
3 The agent approaches the wall from the right, but it is still some distance away. The wall is far away, and the agent is moving in the lower right.
4 The agent is going through the gap in the middle and left, and insert the key at the end of the scene. The agent going through the final gap and inserting the key.
5 The agent approaches the wall from the middle area. We can identify the outline of the agent in the lower middle.
6 The agent going through the gap on the right, and performs policy far from the wall in the right area. The agent is crossing the gap in the bottom right.
7 The agent approaches the wall from the left, but it is still some distance away. The agent is moving in the lower left, and the outline of walls is in the middle of the screen.
We examined a substantial number of video states and corresponding cluster information, and found that the factors determining clusters in FruitBot are the agent’s position on the screen and its relative positioning to walls and gaps. This suggests that the agent has learned critical factors within the environment.
Appendix H Hovering Examples
In figures H.1, and H.2, we present examples of our interactive visualization tool applied to Jumper and FruitBot. This tool is included in the supplementary material, allowing readers to freely explore the semantic distribution of features and gain a better understanding of the semantic clustering properties of DRL.
(a) Overall perspective.
(b) First example point
(c) Second example point
(d) Third example point
Figure H.1: Hover examples in the FDR space of Jumper. We observe a sub-cluster in the FDR space as an example from the overall perspective (a) and the zoomed-in perspective (b), (c), and (d). The agent is standing on the edge of a ledge. Although the scenarios of (b), (c), and (d) are different, the proposed method effectively clusters semantically consistent features together in the FDR space.
(a) Overall perspective.
(b) First example point
(c) Second example point
(d) Third example point
Figure H.2: Hover examples in the FDR space of Fruitbot. We observe a sub-cluster in the FDR space as an example from the overall perspective (a) and the zoomed-in perspective (b), (c), and (d). The agent is standing on the edge of a ledge. Although the scenarios of (b), (c), and (d) are different, the proposed method effectively clusters semantically consistent features together in the FDR space.
Appendix I Human Evaluation Details
I.1 Part 1: Overview and Timeline Introduction
This section provides a detailed description of the evaluation process and presents the timeline for conducting the assessment. The evaluation aims to assess the clarity and interpretability of the semantic clusters in the FruitBot, Jumper, and Ninja games, with the objective of enhancing and quantifying the explainability of the DRL system.
I.1.1 Timeline
Each participant will complete a survey for two game environments (FruitBot, Jumper, or Ninja). They follow the format as detailed:
  • Stage 1: Questionnaire (5 minutes)
    Participants are requested to complete a questionnaire that collects demographic information and gaming-related details. The questionnaire includes sections for gender, age group, education level, occupation, gaming experience, familiarity with evaluating game states, and preferred game genres.
  • Stage 2: Introduction to Game Environment (10 minutes)
    During this stage, participants receive an introduction to the evaluation process. They are informed about the objectives of the assessment and the significance of evaluating the clarity and interpretability of the semantic clusters. They are also given a short description of the game environment (FruitBot, Jumper, or Ninja), and are shown a short gameplay clip to aid with the understanding of the game’s objectives and features.
  • Stage 3: Assessment (50 minutes)
    Following the familiarization period, participants spend 50 minutes assessing the semantic clusters in the game environment. They focus on evaluating the clarity and understandability of the video clips within each semantic cluster. This is done online via a survey.
Total Evaluation Time: 60 minutes.
I.2 Part 2: Questionnaire
This section of the evaluation plan presents the questionnaire that participants are required to complete. The questionnaire consists of the following sections:
I.2.1 Demographic Information
  • Age: Participants indicate their age.
  • Gender: Participants specify their gender as Male, Female, or Other.
  • Education Level: Participants indicate their highest level of education completed, including options such as High school and below, Bachelor’s degree, Master’s degree, and Doctorate and above.
  • Occupation: Participants provide their current occupation, selecting from options such as Student, Employee, Self-employed, or Other.
I.2.2 Gaming-related Information
  • Gaming Experience: Participants indicate their level of gaming experience, choosing from options such as Beginner, Intermediate player, Advanced player, or Professional player.
  • Game Frequency: Participants indicate their frequency of gaming, choosing from options such as Daily, Several times a week, Weekly, Monthly, or others.
  • Experience in Evaluating Game States: Participants assess their experience in evaluating game states, selecting from options such as No experience, Some experience, Moderate experience, or Extensive experience.
  • Preferred Game Genres: Participants specify their preferred game genres, including options such as Role-playing games, Shooting games, Strategy games, Puzzle games, or Other.
I.3 Part 3: Evaluation Questions for Clarity Assessment
In this section, a comprehensive set of questions is provided to assess the clarity and understandability of the semantic clusters. The questions capture participants’ opinions and perceptions using a Likert scale ranging from ‘Strongly Disagree’ to ‘Strongly Agree’. The specific evaluation questions for the clarity assessment include:
  • The clips of each cluster consistently display the same skill being performed.
  • The clips of each cluster match the given skill description.
The two questions are asked each time the participant has been shown a semantic cluster.
I.4 Part 4: Evaluation Questions for Interpretability Assessment
This section outlines the question evaluating the interpretability of the semantic clusters in terms of their usefulness. The question is designed to capture participants’ opinions and perceptions using a Likert scale ranging from "Strongly Disagree" to "Strongly Agree." The specific evaluation question for the interpretability assessment:
  • The identified skills aid in understanding the environment and the AI’s decision-making process.
The above question is asked after the participants have seen all the semantic clusters.
I.5 Part 5: Personnel and Coordination
This section outlines the personnel and coordination aspects of the evaluation. It includes information about evaluator recruitment and compensation. Specifically:
I.5.1 Evaluator Recruitment
15 evaluators are recruited to participate in the evaluation.
I.5.2 Evaluator Compensation
Each evaluator receives \$15 in compensation for their valuable time and contribution to the evaluation process.
By implementing this comprehensive evaluation plan, we gather valuable insights into the clarity and interpretability of the semantic clusters in the FruitBot, Jumper, and Ninja games. The evaluation results provide essential guidance for further quantifying the improved interpretability of DRL models using our proposed method.
I.6 Part 6: Grouping Details
  • Evaluator 1: FruitBot, Jumper
  • Evaluator 2: FruitBot, Jumper
  • Evaluator 3: FruitBot, Jumper
  • Evaluator 4: FruitBot, Jumper
  • Evaluator 5: FruitBot, Jumper
  • Evaluator 6: Jumper, Ninja
  • Evaluator 7: Jumper, Ninja
  • Evaluator 8: Jumper, Ninja
  • Evaluator 9: Jumper, Ninja
  • Evaluator 10: Jumper, Ninja
  • Evaluator 11: Ninja, FruitBot
  • Evaluator 12: Ninja, FruitBot
  • Evaluator 13: Ninja, FruitBot
  • Evaluator 14: Ninja, FruitBot
  • Evaluator 15: Ninja, FruitBot
This grouping plan ensures that each evaluator evaluates two different games, and each game receives a total of 10 evaluations. It allows for comprehensive evaluations of each game and ensures that evaluators have an opportunity to provide feedback on multiple games.
Appendix J Potential Societal Impacts
This paper advances the interpretability of DRL through semantic clustering, with potential applications in safety-critical domains such as autonomous systems and robotics. While primarily contributing to the field of Machine Learning, we encourage responsible application to mitigate potential misuse and do not identify immediate societal or ethical risks requiring specific emphasis.

### v5

@@
 In this paper, we explore semantic clustering properties of deep reinforcement learning (DRL) to improve its interpretability and deepen our understanding of its internal semantic organization.
-In this context, semantic clustering refers to the ability of neural networks to cluster inputs based on their semantic similarity in the internal space.
+In this context, semantic clustering refers to the ability of neural networks to cluster inputs based on their semantic similarity in the feature space.
 We propose a DRL architecture that incorporates a novel semantic clustering module that combines feature dimensionality reduction with online clustering.
@@
 Furthermore, we introduce new analytical methods based on these properties to provide insights into the hierarchical structure of policies and semantic organization within the feature space.
+Our code is available at https://github.com/ualiangzhang/semantic_rl.
 1 Introduction
@@
 In this study, we explore semantic clustering properties to improve the interpretability of DRL models.
-We use the term semantic clustering to refer to the process of grouping information based on underlying semantic similarities in the feature space.
+We use the term semantic clustering to refer to the process of grouping states eliciting similar agent behaviors under comparable environmental contexts (e.g., approaching a target, jumping to a higher platform in Procgen).
 Studying semantic clustering in DRL helps reveal the model’s internal knowledge structure and semantic relationships between states, enhancing the interpretability and transparency of DRL models.
-Although semantic clustering has been thoroughly investigated in natural language processing (NLP) [21, 20] and computer vision (CV) [14, 22, 19, 34], it remains underexplored in DRL, due to the complexity introduced by temporal dynamics and the absence of direct supervised signals.
+Although semantic clustering has been thoroughly investigated in natural language processing (NLP) [4, 5] and computer vision (CV) [6, 7, 8, 9], it remains underexplored in DRL due to the complexity introduced by temporal dynamics and the absence of direct supervised signals.
 The sequential nature of decisions in DRL further complicates the task of capturing evolving semantics.
-Early work introduced external constraints—e.g., bisimulation [9, 33] and contrastive learning [8, 1, 11]—to shape feature spaces conducive to semantic clustering.
+Early work introduced external constraints—e.g., bisimulation [10, 11] and contrastive learning [12, 13, 14, 15]—to shape feature spaces conducive to semantic clustering.
 In contrast, we focus on investigating whether DRL can intrinsically exhibit semantic clustering capabilities.
-Mnih et al.
-[17] and Zahavy et al.
-[32] analyzed the semantic distribution of the DRL feature space for Atari games using t-distributed Stochastic Neighbor Embedding (t-SNE) [27].
-However, these studies are limited in multiple ways: (i) they are limited to a small set of Atari games with fixed scenes, making it difficult to distinguish whether clustering arises from pixel similarity or semantic understanding, (ii) Zahavy et al.
-[32] manually define features for specific games, imposing substantial human effort, and (iii) both studies rely on t-SNE visualization for semantic analysis, which tends to produce unstable results and lacks an automated clustering mechanism.
+[16] and [17] analyzed the semantic distribution of the DRL feature space for Atari games using t-distributed Stochastic Neighbor Embedding (t-SNE) [18].
+However, these studies are limited in multiple ways: (i) they are limited to a small set of Atari games with fixed scenes, making it difficult to distinguish whether clustering arises from pixel similarity or semantic understanding, (ii) [17] manually define features for specific games, imposing substantial human effort, and (iii) both studies rely on t-SNE visualization for semantic analysis, which tends to produce unstable results and lacks an automated clustering mechanism.
 Thus, these approaches require significant manual effort for feature space annotation and analysis, hindering comprehensive semantic analysis and integration into downstream tasks.
@@
   • We comprehensively explore the semantic clustering properties of DRL, advancing the understanding of the black-box decision-making processes.
-Unlike prior work that uses fixed-scene Atari games, we use Procgen  [^1] Detailed environment instructions are available here, in Appendix A of Cobbe et al.
-[4], and in the repository.
+Unlike prior work that uses fixed-scene Atari games, we use Procgen [^1] Detailed environment instructions are available here, in Appendix A of [19], and in the repository.
 [19], which offers rich semantic diversity and dynamic, procedurally generated environments.
-  • We introduce a novel architecture that addresses limitations in previous t-SNE-based semantic analysis methods by providing a more stable and effective approach for studying semantic properties in DRL.
+  • We introduce a novel end-to-end architecture that integrates feature dimensionality reduction with online clustering, overcoming the limitations of prior t-SNE-based analyses and providing a more stable, effective means to study semantic properties in DRL.
   • We present new analysis methods to reveal the internal semantic structure, uncover the hierarchical organization of policies, and identify potential risks in DRL models.
@@
 Similarly, in computer vision, images with similar content are positioned closely in the learned feature space [6, 7, 8, 9].
-Semantic Clustering in DRL Mnih et al.
-[17] and Zahavy et al.
-[32] have previously explored visualizing the DRL feature space using t-SNE.
+Semantic Clustering in DRL [16] and [17] have previously explored visualizing the DRL feature space using t-SNE.
 In these studies, t-SNE visualizations show that features of states with close pixel distances tend to cluster together.
@@
 Interpretability of DRL DRL interpretability research often focuses on video games due to their controlled environments and clear rules, which make analyzing decision-making processes easier.
-PW-Net [10] enhances interpretability by using human-friendly prototypes to explain the model’s decision-making.
-DIGR [30] improves interpretability by generating saliency maps that highlight the most relevant features influencing the agent’s decisions.
+PW-Net [20] uses human-friendly prototypes to explain the model’s decision-making.
+DIGR [21] generates saliency maps tlat highlight the most relevant features influencing the agent’s decisions.
 Concept policy models integrate expert knowledge into multi-agent RL, enabling real-time intervention and interpretation of agent behavior [22].
@@
 Our work explores how DRL models internally organize information, offering new insights into the structure of learned representations.
+VQ-VAE
+VQ-VAE [27] is a family of generative models that combine classic VAE with discrete latent representations through a posterior parameterization.
+Recently, VQ-VAE has been applied to various tasks, including high-resolution image generation [28], video generation [29], and speech coding [30].
+It has also been used in model-based DRL to train transition models [31, 32].
 3 Method
@@
 Background
-The VQ-VAE [26] workflow begins with an encoder network $\hat{E}$ that maps an input $\mathbf{x}$ to a latent representation $\hat{E}(\mathbf{x})$.
+The VQ-VAE workflow begins with an encoder network $\hat{E}$ that maps an input $\mathbf{x}$ to a latent representation $\hat{E}(\mathbf{x})$.
 This representation is then quantized by mapping it to the nearest embedding in a codebook $\left\{\mathbf{e}_{k}|k\in\left\{1,2,\dots,K\right\}\right\}$.
@@
 where $sg$ is a stop-gradient operator and $\beta$ weights the distance reduction between the encoded output $\hat{E}(\mathbf{x})$ and its closest embedding $\mathbf{e}_{k}$.
-In this paper, we modify VQ-VAE to achieve the following objectives: (i) assign features to the nearest VQ embedding for clustering, (ii) seamlessly integrate with DRL training, enabling simultaneous clustering and policy learning, and (iii) enhance clustering and interpretability through joint training with additional losses.
-Further details are provided in subsection 3.2.
+In this paper, we modify VQ-VAE to (i) assign features to the nearest VQ embedding for clustering, (ii) seamlessly integrate with DRL training, enabling simultaneous clustering and policy learning, and (iii) enhance clustering and interpretability through joint training with additional losses.
+Further details are provided in § 3.2.
 Figure 1: Overview of our architecture.
 The upper segment represents the classic DRL training pipeline, while the lower segment introduces the semantic clustering module.
-The Feature Dimensionality Reduction (FDR) net reduces the dimensionality of state features, resulting in FDR features, which the vector quantizer then processes to generate discrete VQ codes (denoted $k$), which represent states associated with clusters, along with the closest VQ embeddings.
+The Feature Dimensionality Reduction (FDR) net reduces the dimensionality of state features, resulting in FDR features, which the vector quantizer then processes to generate discrete VQ codes (denoted kk)—which represent states associated with clusters—along with the closest VQ embeddings.
 Subsequently, $k$ is integrated into the state feature by element-wise addition after being expanded to match the state feature dimensions, enabling conditional policy training that better supports the integration of downstream tasks.
 3.1 Semantic Clustering Module
-To overcome the limitations of previous t-SNE-based semantic analyses (see section 1), we propose a novel semantic clustering module, which includes dimensionality reduction and online clustering.
+To overcome the limitations of previous t-SNE-based semantic analyses (see § 1), we propose a novel semantic clustering module, which includes dimensionality reduction and online clustering.
 Dimensionality Reduction Given the complexity of states in most DRL applications, their features are often high-dimensional.
-For example, Mnih et al.
-[17] use a 512-dimensional feature vector when training DQN on Atari games.
-Clustering high-dimensional features is challenging due to "the curse of dimensionality" [3, 2].
+For example, [16] use a 512-dimensional feature vector when training DQN on Atari games.
+Clustering high-dimensional features is challenging due to the curse of dimensionality [33, 34].
 To mitigate these issues, we reduce feature dimensionality before clustering, resulting in more robust clustering outcomes.
@@
 This network remaps high-dimensional features to 2D using policy training data for online training, ensuring stable and efficient mappings after training.
-The FDR network’s loss function is designed to preserve the consistency of distance relationships between high-dimensional and 2D feature spaces, measured by pairwise similarities as described in subsection 3.2.
-Online Clustering To reduce the need for extensive manual annotation from previous work and to facilitate the integration of downstream tasks, we propose using online clustering to segment the feature space and assist in semantic analysis.
-We modify the VQ-VAE framework to achieve online clustering, as detailed in subsection 3.2.
+The FDR network’s loss function is designed to preserve the consistency of distance relationships between high-dimensional and 2D feature spaces, measured by pairwise similarities as described in § 3.2.
+Online Clustering
+t-SNE–based analyses (e.g., [16, 17]) require per-state inspection and manual grouping because the plots lack clear cluster boundaries and often split semantically similar states across disconnected regions, making human curation time-consuming.
+To reduce such extensive annotation and facilitate downstream integration, we introduce an online clustering approach—implemented via a modified VQ-VAE—that automatically segments the feature space and supports semantic analysis.
+This removes manual grouping: annotators instead watch a few short clips per discovered cluster and provide a semantic summary (see Table 1), typically within $\approx$15 minutes per environment.
+Details of the modified VQ-VAE design are provided in § 3.2.
 3.2 Loss Function Design
 The loss function for our proposed framework is given by
-$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{DRL}}+f_{\text{control}}\left(w_{\text{FDR}}\mathcal{L}_{\text{FDR}}+w_{\text{VQ-VAE}}\mathcal{L}^{\prime}_{\text{VQ-VAE}}\right).$ (2)
+$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{DRL}}+\lambda_{\text{ctrl}}\left(w_{\text{FDR}}\mathcal{L}_{\text{FDR}}+w_{\text{VQ-VAE}}\mathcal{L}^{\prime}_{\text{VQ-VAE}}\right).$ (2)
 The DRL loss function $\mathcal{L}_{\text{DRL}}$ comes from the original DRL model.
 $w_{\text{FDR}}$ and $w_{\text{VQ-VAE}}$ are the weights of the FDR loss ($\mathcal{L}_{\text{FDR}}$) and the modified VQ-VAE loss ($\mathcal{L}^{\prime}_{\text{VQ-VAE}}$), respectively.
-$f_{\text{control}}$ represents the control factor.
+$\lambda_{\text{ctrl}}$ represents the control factor.
 We explain each of these components below.
@@
 We use the Student’s t-distribution for pairwise similarities as it captures nonlinear structures and efficiently measures pairwise relative positions of features within a batch without requiring the entire feature set, making it ideal for online clustering.
-The pairwise similarities of state features $p_{ij}$ is given by
+The pairwise similarities of state features $p_{ij}$ are given by
 $p_{ij}=\frac{d(i,j)}{\sum_{k\neq l}d(k,l)},\quad\text{where }d(m,n)=\left(1+\frac{\left\|f(\mathbf{s}_{m})-f(\mathbf{s}_{n})\right\|^{2}}{\alpha}\right)^{-\frac{\alpha+1}{2}}.$ (3)
-where $f$ is the feature extractor, $\mathbf{s}_{i}$ is the $i^{\text{th}}$ state in a batch, and $\alpha$ is the degree of freedom.
+Here, $f$ is the feature extractor, $\mathbf{s}_{i}$ is the $i^{\text{th}}$ state in a batch, $\alpha$ is the Student’s-t degrees-of-freedom parameter, and $\|\cdot\|$ is the $\ell_{2}$ norm.
 The pairwise similarities for FDR features, $q_{ij}$ are computed using the same expression as (3), but with $f$ replaced by $g\circ f$, where $g$ is the FDR net.
-In contrast to other deep clustering studies, e.g., Xie et al.
-[29] and Li et al.
-[12], the same degree of freedom $\alpha$ is selected for both high- and low-dimensional similarities, ensuring that the original distance relationship between features is maintained in the low-dimensional space.
+In contrast to other deep clustering studies, e.g., [36] and [37], the same degree of freedom $\alpha$ is selected for both high- and low-dimensional similarities, ensuring that the original distance relationship between features is maintained in the low-dimensional space.
 The FDR loss is given by
 $\begin{split}\mathcal{L}_{\text{FDR}}&=-\sum_{i}\sum_{j}p_{ij}\log\left(q_{ij}\right).\end{split}$ (4)
+Minimizing $\mathcal{L}_{\text{FDR}}$ encourages the low‑dimensional mapping to preserve the pairwise neighborhood structure of the high‑dimensional features.
 Modified VQ-VAE Loss To perform clustering, we use the second term of $\mathcal{L}_{\text{VQ-VAE}}$ from (1), which moves VQ embeddings closer to neighboring FDR features.
@@
 where $\mathbf{e}_{k}$ is the closest embedding in the codebook to the FDR feature $g(f(\mathbf{s}))$.
-Control Factor Since effective semantic clustering relies on a clear and distinguishable semantic distribution that is often difficult to achieve in the early stages of training, we propose an adaptive control factor ($f_{\text{control}}$) strategy updated according to training performance (see Appendix A).
-Improved Clustering Our loss design not only achieves dimensionality reduction and clustering but also enhances clustering properties, making the states within each cluster more compact and the cluster boundaries more separable.
-This improvement is crucial for clearly distinguishing the semantics of states at the cluster boundaries, further enhancing the model’s interpretability.
-We demonstrate the improved clustering in section 4 and provide more evidence of this enhanced clustering property and the intrinsic nature of semantic clustering in DRL in subsection C.2.
-This enhanced clustering arises from the fact that although the stop-gradient operation ($sg$) in $\mathcal{L}^{\prime}_{\text{VQ-VAE}}$ prevents FDR features from moving towards the closest VQ embeddings, when the FDR features become denser under the influence of $\mathcal{L}_{\text{FDR}}$, it indirectly reduces $\mathcal{L}^{\prime}_{\text{VQ-VAE}}$ and, consequently, $\mathcal{L}_{\text{total}}$.
-In other words, training with both $\mathcal{L}_{\text{FDR}}$ and $\mathcal{L}^{\prime}_{\text{VQ-VAE}}$ decreases the sum of distances between FDR features and their corresponding VQ embeddings within the same cluster.
-Additionally, since $\mathcal{L}_{\text{FDR}}$ aims to bring $p$ and $q$ closer, the improved cluster density in the FDR feature space is also transferred to the state feature space.
+Control Factor Since effective semantic clustering relies on a clear and distinguishable semantic distribution that is often difficult to achieve in the early stages of training, we propose an adaptive control factor ($\lambda_{\text{ctrl}}$) strategy updated according to training performance (see Appendix A).
+Improved Clustering Our loss design not only achieves dimensionality reduction and clustering but also enhances clustering properties, making the states within each cluster more compact (smaller intra‑cluster distances) and the cluster boundaries more separable.
+This is crucial for clearly distinguishing the semantics of states at the cluster boundaries, further enhancing the model’s interpretability.
+Because of the stop-gradient in $\mathcal{L}^{\prime}_{\text{VQ-VAE}}$, it does not directly pull FDR features toward their nearest codebook embeddings.
+However, when the FDR features become denser during joint training with $\mathcal{L}_{\text{FDR}}$, $\mathcal{L}^{\prime}_{\text{VQ-VAE}}$—and thus $\mathcal{L}_{\text{total}}$—decreases, yielding tighter clusters.
+Moreover, since $\mathcal{L}_{\text{FDR}}$ aligns the affinity matrices $p$ and $q$, this densification in the low-dimensional FDR space is reflected and propagates into the high-dimensional state features.
+We demonstrate the improved clustering in § 4 and provide more evidence of this enhanced clustering property and the intrinsic nature of semantic clustering in DRL in § C.2.
 1
 Input: PPO network parameters $\theta$, FDR network parameters $\phi$, SCM hyperparameters, and PPO hyperparameters such as value loss weight $w_{\text{value}}$, entropy loss weight $w_{\text{entropy}}$.
-2 for each training iteration $i=1,2,\dots$ do
+2
+31exfor each training iteration $i=1,2,\dots$ do
   Collect NN trajectories $\mathcal{D}_{i}=\{\tau_{1},\dots,\tau_{N}\}$ using policy $\pi_{\theta}$;
@@
           // Compute value outputs
-6
-       $\mathcal{L}_{\text{PPO}}\leftarrow\mathcal{L}_{\text{policy}}+w_{\text{value}}\cdot\mathcal{L}_{\text{value}}+w_{\text{entropy}}\cdot\mathcal{L}_{\text{entropy}}$;
+7
+       $\mathcal{L}_{\text{PPO}}\leftarrow\mathcal{L}_{\text{policy}}+w_{\text{value}}\mathcal{L}_{\text{value}}+w_{\text{entropy}}\mathcal{L}_{\text{entropy}}$;
        // PPO loss
-       $\mathcal{L}_{\text{SCM}}\leftarrow w_{\text{FDR}}\cdot\mathcal{L}_{\text{FDR}}+w_{\text{VQ-VAE}}\cdot\mathcal{L}^{\prime}_{\text{VQ-VAE}}$;
+       $\mathcal{L}_{\text{SCM}}\leftarrow w_{\text{FDR}}\mathcal{L}_{\text{FDR}}+w_{\text{VQ-VAE}}\mathcal{L}^{\prime}_{\text{VQ-VAE}}$;
        // SCM loss
-       $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{PPO}}+f_{\text{control}}\cdot\mathcal{L}_{\text{SCM}}$;
+       $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{PPO}}+\lambda_{\text{ctrl}}\mathcal{L}_{\text{SCM}}$;
        // Total loss
@@
        // Parameter update
-7
 8
 9
+10
 Algorithm 1 PPO with Semantic Clustering Module (SCM)
-Advantages of Online Training Online training offers several key advantages: (i) It enhances clustering by incorporating the training of $\mathcal{L}_{\text{total}}$.
-(ii) Training the VQ code $k$ with a latent-conditioned policy $\pi\left(a|\mathbf{s},k\right)$ (where $a$ is the action) supports the extension to downstream tasks, such as macro action selection in hierarchical learning.
-(iii) Online training improves memory efficiency by eliminating the need to store a large number of states during model training.
+Advantages of Online Training Online training offers several advantages: (i) it enhances clustering by incorporating the training of $\mathcal{L}_{\text{total}}$, (ii) training the VQ code $k$ with a latent-conditioned policy $\pi\left(a|\mathbf{s},k\right)$ (where $a$ is the action) supports extension to downstream tasks, such as macro action selection in hierarchical learning, and (iii) it improves memory efficiency by eliminating the need to store a large number of states during model training.
 Training Process The training process of our framework builds upon the structure of the original DRL algorithm while incorporating the semantic clustering module (SCM) by using (2) for total loss calculation.
@@
 In this work, we primarily study the intrinsic characteristics and black-box decision-making of DRL, and address the instability of t-SNE visualizations used in prior studies.
-Therefore, this section aims to: (i) compare t-SNE to validate the stability and effectiveness of the proposed clustering method; (ii) assess the semantic clustering properties of DRL to improve interpretability; and (iii) introduce new methods to analyze policies and internal model characteristics, identifying issues in DRL decision-making.
-For performance analysis, please refer to subsection C.1 and Appendix D, which demonstrate that the integration of our module has minimal impact on performance.
+Therefore, this section aims to: (i) compare t-SNE to validate the stability and effectiveness of the proposed clustering method, (ii) assess the semantic clustering properties of DRL to improve interpretability, and (iii) introduce new methods to analyze policies and internal model characteristics, identifying issues in DRL decision-making.
+The integration of our module has minimal impact on performance (see § C.1 and Appendix D).
 4.1 Clustering Effectiveness Evaluation
@@
 We use a trained model to collect states, where the agent selects actions randomly with a probability of 0.2 and follows the trained policy with a probability of 0.8 to ensure diverse state coverage.
-States are sampled with a probability of 0.8, and 64 parallel environments collect states over 500 steps, resulting in $\sim$25,000 states for visualization.
+States are sampled with a probability of 0.8, and 64 parallel environments collect states over 500 steps, resulting in $\approx$25,000 states for visualization.
+Note that the cluster colors (indices) in the t-SNE plots are assigned by our method and are used solely to facilitate comparison of spatial relationships.
 Cluster Separation and Improved Clustering The t-SNE visualization of PPO (2(a)), spreads features across the space without forming clear clusters, limiting its utility for clustering analysis and requiring detailed manual examination of certain areas, as in previous studies.
@@
 The visualization in the FDR space (2(d)), displays clear and separate complete clusters, which are identified by VQ codes.
+§ C.2 presents a stop-gradient ablation that disables our proposed module while keeping the backbone and training protocol fixed; visualizations show fuzzier, less separable clusters than the full model, further supporting our gains in sharpness and coherence.
 Sensitivity to Number of States Our method’s stability is showcased in Figures 2(c) and 2(e), where the number of processed states is reduced by 50%.
@@
 Sensitivity to Random Seed While the t-SNE representation is sensitive to randomness, as demonstrated by the significant difference between Figures 2(c) and 2(f), the FDR space’s mapping remains unchanged even when the random seed is altered, maintaining the distribution in 2(e).
-t-SNE’s randomness primarily stems from its random initialization and non-convex optimization process, leading to different visualizations with different random seeds.
+t-SNE’s randomness primarily stems from its random initialization and non-convex optimization process, leading to significantly different visualizations with different random seeds.
 In contrast, our model produces a stable feature mapping after training, which does not vary with random seeds.
@@
 Human Evaluation In addition to qualitatively analyzing the mean images, we hired 15 human evaluators to validate the semantic clustering properties.
+Evaluators were adults (18+), native or highly proficient English speakers, with basic video game experience and a brief training session.
 Specifically, video sequences from each episode are segmented into multiple clips based on the cluster each frame belongs to, and these clips are grouped by cluster for evaluators to review.
 Each evaluator watched these grouped clips and responded to three interpretability-related statements for two out of a set of three games (Jumper, Fruitbot, and Ninja).
-The response for each question was chosen from a five-point Likert scale with the following options: Strongly Disagree (1), Disagree (2), Neutral (3), Agree (4), and Strongly Agree (5).
-Further details on the human evaluation procedure are provided in Appendix I.
-The statements and the results of the human evaluation are provided in subsection 4.2.
+The response for each question was chosen from a five-point Likert scale with the options: Strongly Disagree (1), Disagree (2), Neutral (3), Agree (4), and Strongly Agree (5).
+Further details on the evaluation procedure are provided in Appendix I.
+The statements and the results of the human evaluation are provided in § 4.2.
 The mean scores for all statement-environment combinations are greater than 4, with the exception of statements 1 and 3 for FruitBot, for which the lower bounds on the mean set by the standard error of the mean (SEM) are 3.99 and 3.92 respectively.
@@
 Figure 5: Hover examples in the FDR space of Ninja.
-We observe a sub-cluster in the FDR space as an example from the overall perspective (a) and the zoomed-in perspective (b), (c), and (d).
+We observe a sub-cluster in the FDR space as an example from an zoomed-out perspective (a) and zoomed-in perspectives (b), (c), and (d).
 The agent is standing on the edge of a ledge.
@@
 The tool supports ‘statically’ analyzing the semantic distribution of models—specifically, (i) when the mouse cursor hovers over a specific feature point, the corresponding state image is displayed, and (ii) the tool includes a zooming functionality to observe the semantic distribution of features in detail within clusters.
-Figure 6: Policy analysis examples in Ninja.
+Figure 6: Policy analysis examples in Ninja, showing states assigned to different clusters.
 In addition, we propose a more ‘dynamic’ analysis method—the VQ code enables us to determine the cluster to which the current state belongs, which allows for the semantic segmentation of episodes, as exemplified in Figure 4.
@@
 We present policy analysis examples in Figure 6, leveraging clustering results from our method.
-Figures 6(a) and 6(b) illustrate consecutive states assigned to clusters 5 and 7, respectively.
-These examples show that the decision to approach the mushroom depends not on the mushroom’s presence but on detecting a right-side wall.
-This finding is confirmed in 6(c), another state from cluster 7 without any mushroom present, where the agent repeatedly moves along the ledge, incorrectly perceiving conditions for behaviors in cluster 7.
+Figures 6(a) and 6(b) show consecutive states assigned to clusters 5 and 7, respectively (see Figure 3 and Table 1).
+In 6(a), the right-side wall is absent and the agent walks along the ledge (cluster 5); in 6(b), a right-side wall appears and the agent transitions to cluster 7, approaching the mushroom.
+Because the mushroom is visible in both frames, this cluster change is driven by detecting the right-side wall rather than by the mushroom’s presence.
+This finding is confirmed in 6(c), which shows another state from cluster 7 without any mushroom present, where the agent continues along the ledge, incorrectly perceiving the conditions for cluster-7 behavior.
 Figures 6(d) and 6(e) depict states within the same episode: in 6(d), the agent initially plans to jump onto a lower ledge (cluster 5), but upon seeing a higher, safer ledge in 6(e), it shifts its strategy accordingly (cluster 3).
@@
 More robust clustering methods may be needed to improve stability.
-Second, as the method is unsupervised, selecting an appropriate number of clusters is crucial—too few reduce clarity, while too many cause semantic fragmentation.
-We used eight clusters to balance interpretability and granularity, but future work could explore adaptive techniques that adjust the number of clusters based on task complexity.
+Second, as the method is unsupervised, selecting an appropriate number of clusters is crucial—too few clusters reduces clarity, while too many clusters causes semantic fragmentation.
+We used eight clusters to balance interpretability and granularity, but future work could explore adaptive techniques that adjust the number of clusters based on task complexity, the elbow method, silhouette-score optimization, etc.
 For further analysis of the impact of cluster numbers on performance and interpretability, see Appendix D.
 Furthermore, policy interpretations are manually described.
-In future work, we aim to automate behavior summarization and explanation.
+In future work, we aim to automate behavior summarization and explanation (e.g., using GPT-4V).
+In addition, since the FDR module optimizes pairwise similarities, a natural extension is to replace the current affinity with alternative measures (e.g., cosine similarity or bisimulation metrics).
 Lastly, we plan to extend this method to other DRL algorithms, benchmarks, and settings.
@@
 This dynamic clustering exploits regularities in the environment, offering a unique approach compared to the static clustering observations in NLP and CV.
+Acknowledgments
+Research was sponsored by the Army Research Office and was accomplished under awards W911NF-20-1-0002 and W911NF-24-2-0034.
+The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S.
+Government.
+The U.S.
+Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
+We gratefully acknowledge the anonymous reviewers and the program committee for their thoughtful feedback and constructive suggestions, which helped improve this work.
+We also thank Huy Le and Robert Lopez for their valuable assistance with the human evaluation.
 Appendix Overview
@@
 Each runs takes around six hours.
-In Equation 2 of the main paper, $w_{\text{FDR}}$ and $w_{\text{VQ-VAE}}$ are 500 and 1, respectively. $f_{\text{control}}$ is updated every 50 iterations according to the following expression:
-$f_{\text{control}}=\min\left(\frac{s_{\text{mean}}}{0.8\cdot s_{\text{highest}}},1\right),$ (A.1)
+In Equation 2 of the main paper, $w_{\text{FDR}}$ and $w_{\text{VQ-VAE}}$ are 500 and 1, respectively. $\lambda_{\text{ctrl}}$ is updated every 50 iterations according to the following expression:
+$\lambda_{\text{ctrl}}=\min\left(\frac{s_{\text{mean}}}{0.8\cdot s_{\text{highest}}},1\right),$ (A.1)
 where $s_{\text{mean}}$ is the mean score of the last 100 episodes in training, and $s_{\text{highest}}$ is the highest score of the environment.
@@
 The number of embeddings in the vector quantizer was determined by ensuring that each cluster maintained a singular semantic interpretation.
+During hyperparameter tuning, we found that performance is primarily influenced by $w_{\text{FDR}}$, $w_{\text{VQ-VAE}}$, and $\lambda_{\text{ctrl}}$, and is more robust to the number of VQ embeddings and the degrees of freedom in Equation 3.
 Appendix B Theoretical Analysis of Loss Design
@@
 Bars show mean values; error bars denote the standard error of the mean.
-To quantify how well various dimensionality‑reduction (DR) techniques preserve the semantic clusters discovered by our model, we project the same set of high‑dimensional state features—collected following the procedure described in subsection 4.1—using five popular DR baselines: UMAP, Isomap, PCA, MDS, and t‑SNE.
+To quantify how well various dimensionality‑reduction (DR) techniques preserve the semantic clusters discovered by our model, we project the same set of high‑dimensional state features—collected following the procedure described in § 4.1—using five popular DR baselines: UMAP, Isomap, PCA, MDS, and t‑SNE.
 These metrics were selected to evaluate both the compactness within clusters and the separation between clusters in the reduced feature space.
@@
 Appendix F Semantic Formation in Clusters
-To provide insight into the generation of semantic clustering in the feature space, we conducted the statistical analysis as follows.
-Initially, we collected 50,000 states using the trained model and computed the mean image for each cluster.
-We then calculated the mean and standard deviation of the pixel distances between the states in each cluster and their corresponding mean image.
-These measurements were averaged across the eight clusters.
-Additionally, we evaluated the probability of cluster transitions in these 50,000 states.
-We repeated the same statistical analysis on the 50,000 states using the untrained model, and the results are presented in Table 4.
-In the trained models, the pixel differences between the states within clusters increased, indicating the presence of states with similar semantics but larger pixel distances.
-Furthermore, the probability of cluster transitions decreased as it approached the required transition in the scene.
-Table 4: Comparison metrics of the semantically trained model verse the raw model on various environments.
-Model Cluster transition probability Pixel distance mean (Std. Dev.)
+To analyze how semantic clusters form in the feature space, we sample 50,000 states per environment and compute: (i) the per-cluster mean image and the mean (± std) pixel distance from each state to its cluster mean (averaged over K=8K{=}8 clusters), and (ii) the probability of cluster transitions along trajectories.
+We evaluate three models: Trained (ours), Stop-Grad (the ablation in § C.2 that removes the effect of our modules by stopping gradients), and Raw (untrained).
+Results are shown in Table 4.
+Compared to Raw, both Trained and Stop-Grad reduce transition probability, indicating policy-induced structure.
+However, Stop-Grad lacks fully distinct boundaries: it exhibits higher transition probability and lower intra-cluster pixel distance than Trained, whereas our method achieves the lowest transition probability and the highest intra-cluster pixel distance.
+Table 4: Cluster transition probability and intra-cluster pixel distance (mean with std) over 50k states.
+Environment Model Cluster transition probability Pixel distance mean (Std. Dev.)
 FruitBot Trained 0.1081 100.00 (71.33)
+FruitBot Stop-Grad 0.1520 93.21 (68.09)
 FruitBot Raw 0.2834 77.10 (49.94)
 Jumper Trained 0.2224 110.29 (62.29)
+Jumper Stop-Grad 0.3015 108.38 (61.22)
 Jumper Raw 0.5829 104.57 (58.31)
-Ninja Trained 0.2680 141.50 (67.88)
+Ninja Trained 0.2680 141.57 (67.88)
+Ninja Stop-Grad 0.2705 132.46 (66.12)
 Ninja Raw 0.2712 87.61 (62.43)
+We further assess temporal coherence with two episode-level metrics: (i) Episode Cluster Entropy (ECE)—the entropy of each episode’s cluster distribution (lower is better, indicating more focused semantic grouping), and (ii) Temporal Cluster Agreement $\text{TCA@}k$—the fraction of frame pairs at lag $k$ assigned to the same cluster (higher is better, indicating smoother, more stable semantics).
+As summarized in Table 5, our method consistently achieves lower ECE and higher TCA@3/6 than Stop-Grad across all three games.
+Note that Jumper episodes are shorter, yielding fewer clusters per episode and thus lower ECE values overall.
+Table 5: Episode- and frame-wise metrics (averaged over episodes).
+Environment Model ECE TCA@3 TCA@6
+FruitBot Stop-Grad 1.9068 0.7050 0.5309
+FruitBot Ours 1.8886 0.7302 0.5502
+Jumper Stop-Grad 0.6987 0.7040 0.6715
+Jumper Ours 0.5572 0.7911 0.7089
+Ninja Stop-Grad 1.5505 0.5798 0.3923
+Ninja Ours 1.3086 0.6787 0.4985
 Appendix G More Examples and Mean Images in the FDR Space
@@
 While primarily contributing to the field of Machine Learning, we encourage responsible application to mitigate potential misuse and do not identify immediate societal or ethical risks requiring specific emphasis.
+NeurIPS Paper Checklist
+The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
+Do not remove the checklist: The papers not including the checklist will be desk rejected.
+The checklist should follow the references and follow the (optional) supplemental material.
+The checklist does NOT count towards the page limit.
+Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
+  • You should answer [Yes] , [No] , or [N/A] .
+  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
+  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
+The checklist answers are an integral part of your paper submission.
+They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
+You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
+The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation.
+While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used").
+In general, answering "[No] " or "[N/A] " is not grounds for rejection.
+While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate.
+All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix.
+If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
+IMPORTANT, please:
+  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
+  • Keep the checklist subsection headings, questions/answers and guidelines below.
+  • Do not modify the questions and only use the provided macros for your answers.
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The abstract and the introduction precisely enumerate three key contributions.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: A dedicated paragraph (§5) details the limitations.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: Appendix B states all assumptions and supplies full proofs.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: §4 and App. A list environment names, seeds, architecture, hyper‑parameters (learning rates, loss weights), compute budget, and provide the code in the supplementary materials.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: The supplementary zip file contains anonymized source code, configuration files, and scripts that fetch the public Procgen dataset.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: Training details (optimiser, batch‑size, PPO clip, etc.) appear in App. A; evaluation follows the Procgen protocol exactly.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: All learning curves show mean and error bars.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: App. A states that each run uses a single Tesla V100 (32 GB) for 6 h; total GPU‑days are reported.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The work uses only synthetic game data, releases code under MIT license, and poses no identifiable privacy or safety risk.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: Appendix J discusses the impacts of this study.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The released artifacts are small game agents that carry no foreseeable dual-use risk.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We cite the Procgen benchmark [19] and note its MIT license in App. A.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: No new datasets or models are released beyond trained weights already covered by the code‑release answer.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: Our user study surveyed adult volunteers on AI models online without studying human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: Our user study surveyed adult volunteers online, posed minimal risk, and was exempt under our institution’s IRB policy.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [No]
+  79. Justification: No large language models are used in the proposed methodology.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.